OpenStudy (anonymous):

Let u = <-6, 1>, v = <-5, 2>. Find -4u + 2v.

OpenStudy (anonymous):

@AloneS

OpenStudy (neonumbrella5115):

What do you think the answer is?

OpenStudy (anonymous):

@neonumbrella5115 i have no idea

OpenStudy (wilsondanielle):

If I'm interpreting the question correctly, it's telling you that the value of u is somewhere between -6 and 1 and the value of v is somewhere between -5 and 2. Because you don't have an exact value, you will need to express your answer as a range. Start by calculating the final answer with the lower values, and then calculate the answer with the higher values. This will be your range within which the answer will lie. eg: Lower values Lowest range of u : -6 Lowest range of v : -5 -4u+2v = -4*-6 + 2*-5 = 24-10 = 14 Repeat this process with the larger values, and express your final answer as a range.